# Resource-constrained Fairness

### v1

@@ Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources mean that this is not possible. Existing tools for fair machine learning ignore these key constraints, with the majority of methods disregarding any finite resource limitations under which decisions are made. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within this framework. We demonstrate that the level of available resources significantly influences this cost, a factor that has been overlooked in previous evaluations.
Keywords Fairness  $\cdot$ Constrained resources  $\cdot$ Responsible AI
@@ 1 Introduction
However, limited resources are not taken into account in current discussions in fair machine learning. Goethals et al. (2024) show that the number of selected individuals can vary substantially after deploying different bias mitigation methods, including learning fair representations (Zemel et al., 2013) and adversarial debiasing (Zhang et al., 2018). A couple of notable exceptions stand out: Kwegyir-Aggrey et al. (2023) discussed how practitioners need to adapt the threshold to ensure the outcomes meet their domain-specific needs, while Corbett-Davies and Goel (2018) explored constraining the number of selected instances to satisfy a certain budget.
In the standard setting of an unconstrained budget, Mittelstadt et al. (2023) posit that diminishing the performance of privileged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling-down’. We agree with this viewpoint but argue that within contexts constrained by limited resources, some rebalancing is necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources to ensure equality could also be termed ‘leveling down’.
@@
  • We reformulate the problem of fair machine learning as a resource-constrained problem, where we treat the positive decisions as resources to allocate among different groups.
  • We provide a theoretical connection to leveling up, demonstrating that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We quantify the cost of fairness in this resource-constrained framework. Our formulation allows us to examine the factors that affect this cost of fairness. We study the impact of factors inherent to the dataset, including overall uncertainty and per group base rate disparity, uncertainty, and size, alongside context-related factors such as the resource level and enforced fairness metric. We find that the selected positive decision rate has a substantial impact on this cost, an aspect that is not taken into account in previous evaluations. This failure to account for real-world constraints on fairness may be a contributing factor to the field’s lack of prior cases of enforcing fairness on production models, which contrasts starkly with the abundance of studies using fairness metrics solely for model testing and debugging.
@@ 2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$. Let $\mathrm{H}_{\cal D}[c_{w}]$be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\cal D$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
$\mathrm{H}_{\cal D}[c_{w}]=\mathrm{E}_{x\in\cal D}H(y_{x},c_{w}(x))$ (1)
Such measures of a harm $H$ might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases that go unflagged for follow-up treatment).
When we are concerned about fairness in classification, we usually want to measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate (Verma and Rubin, 2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic-loss) while ensuring that the harm is the same per group.
This means that we are searching for a solution to the following problem (where $\cal G$ is a partitioning of the distribution into groups with respect to a particular protected attribute):
$\min_{w}\,\mathrm{E}_{x\in\cal D}\ell(y_{x},c_{w}(x))\text{ such that }$ (2)
$\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\cal G$
@@ 2.1 Leveling down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level-down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in some groups receiving a lower rate of cancer detection than they would otherwise). Note that this process of enforcing fairness can alter the overall selection rate of the model (Goethals et al., 2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $k$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\cal D}\ell(y_{x},c_{w}(x))\text{ such that }\mathrm{H}_{g}[c_{w}]\leq k\,\forall g\in\cal G$ (3)
This is the same as saying that in the worst case, the harm should be below k, for any group. The challenge now becomes, “How should $k$ be set?” When deciding to deploy a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can be used to select $k$ per group.
@@
There are multiple possible answers here. One is a matter of personal utility, that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy, (Menon and Williamson, 2018) and fairness on the ground, which measures the effect of the actual implementation of fairness principles in real-world situations (Bakalar et al., 2021). Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, real-world scenarios often lack the necessary resources to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from a testing to production environment.
@@ 2.2 Constrained resources
We formalize the notion of resources, as the number of instances predicted as positive by the machine learning model. This term can be used interchangeably with capacity, or with selection rate or positive decision rate when speak about the proportion of positively predicted instances. We define $\mathrm{R}_{\cal D}[c_{w}]$to be the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\cal D$.
@@
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\text{ such that }\mathrm{R}_{\cal D}[c_{w}]\leq R$ (4)
where $\mathrm{R}_{\cal D}[c_{w}]$is a weighted sum of group-specific rates $\mathrm{R}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{R}_{\cal D}[c_{w}]$strictly increasing with respect to $\mathrm{R}_{g}[c_{w}]$.
@@
Define $W$ as the set of worst-off groups under the harm $H$, such that:
@@
The optimum must occur when $\mathrm{R}_{\cal D}[c_{w}]=R$. If not, one could distribute the remaining resources such that the rates $\mathrm{R}_{g}[c_{w}]$ for the worst-off groups increase without exceeding R, thus reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{R}_{g}[c_{w}]$).
Furthermore, the optimum is reached when the harm levels are as equal as possible across all groups (so all groups are in $W$). If this were not the case, then there would be at least one group not in $W$, that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{R}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{R}_{\cal D}[c_{w}]\leq R$.
This proof also holds when the harm considered is strictly increasing with respect to $\mathrm{R}_{g}[c_{w}]$ (e.g., a harm corresponding to $1-\text{precision}$). In this case, the optimal solution is trivial with an overall selection rate of 0, and equality will be preserved. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some R. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. By revisiting the previous proof, we find that equality must hold.
Hence, we find that equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below). ∎
@@ 2.3 Cost of fairness
With this in mind, we consider the following question:
What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
We have shown that under constrained resources, this leads to the same results as enforcing fairness, i.e. equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
$\mathrm{H}_{\cal D}[c^{f}_{w}]-\mathrm{H}_{\cal D}[c^{o}_{w}]$
@@
In some cases, we might also be interested in other performance metrics than the fairness metric we enforce. If we want to quantify the loss in precision, we could ask:
What is the change in precision if we minimize the harm of the worst-off group?
This represents a common scenario. In a medical context it corresponds to the question what proportion of healthy patients instead of sick will we see if we increase test sensitivity for the worst off groups? We measure this cost and the factors that contribute to it in our results.
Both questions can also lead to a follow-up question:
What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we minimise the harm of the worst off-group?
@@
The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increases the resources available, and target these resources at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources. The number of available resources can be an alterable parameter that should be taken into account.
3 Materials and methods
3.1 Materials
@@
We use several real world datasets that are common in the domain of fair machine learning (Le Quy et al., 2022). Besides these often used datasets from the tabular domain, we also extend our findings to two datasets (CelebA Liu et al. (2015) and Fitzpatrick17k Groh et al. (2021)) from computer vision. The Adult Income dataset comprises information from the 1994 census data, focusing on whether an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. (2020) and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We prepossess the data and binarize the labels following the approach of Zong et al. (2022). We represent each dataset and specify the protected and target attributes Table 1. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $s$ representing the protected group and $ns$ the privileged group, then it can be mathematically expressed as:
@@
3.2 Methods
3.2.1 Machine learning model
For the tabular datasets, we use the eXtreme Gradient Boosting (XGBoost) model, a gradient boosting framework optimized for both speed and performance Chen and Guestrin (2016). We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. (2016) pretrained on ImageNet Deng et al. (2009) for feature extraction.

### v3

@@ Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources mean that this is not possible. When deploying machine learning systems, these resource constraints are simply enforced by varying the threshold of a classifier. However, these finite resource limitations are disregarded by most existing tools for fair machine learning, which do not allow the specification of resource limitations and do not remain fair when varying thresholds. This makes them ill-suited for real-world deployment. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within this framework. We demonstrate that the level of available resources significantly influences this cost, a factor overlooked in previous evaluations.
Keywords Algorithmic Fairness  $\cdot$ Responsible AI in practice
@@ 1 Introduction
Figure 1: Illustration of how the different bias mitigation methods enforce fairness at a specific selection rate (or threshold) but break when we consider the full range of possible selection rates (Adult Income dataset) . [^2] The icons represent the labels of the bias mitigation methods at one specific threshold. The dashed line represent the output for the other thresholds, by using the prediction scores generated by that method. Note that this approach is not possible for Fairlearn as they do not return a prediction score.
However, limited resources are not taken into account in most discussions in fair machine learning. Kwegyir-Aggrey et al. (2023) also highlight how practitioners typically need to adapt the threshold to ensure that outcomes meet their domain-specific needs, while corbett2023measure argue that the trade-offs in fair decision-making are most acute when there are constrained resources. This failure to account for real-world constraints on fairness may be a contributing factor to the field’s lack of prior cases of enforcing fairness on production models, which contrasts starkly with the abundance of studies using fairness metrics solely for model testing and debugging.
Figure 2 illustrates how demographic parity, one of the most common notions of fairness, varies with selection rate. Many of the currently popular bias mitigation methods, enforced using a range of tools from AIF360 Bellamy et al. (2018) and Fairlearn Bird et al. (2020), aim to ensure fairness at a specific threshold. However, as the thresholds vary, fairness oscillates wildly, revealing that these methods may break when considering the whole range of possible selection rates.
In the standard setting of an unconstrained budget, Mittelstadt et al. (2023) posit that increasing harm to privileged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling down’. We agree with this viewpoint but argue that within contexts constrained by limited resources, some rebalancing is necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources to ensure equality could also be termed ‘leveling down’. .
@@
  • We reformulate the problem of fair machine learning as a resource-constrained problem, where we treat the positive decisions as resources to allocate among different groups. We propose an algorithm that enforces fairness under resource limitations.
  • We provide a connection to leveling up, demonstrating that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We quantify the cost of fairness in this resource-constrained framework. We study the impact of factors inherent to the dataset. We find that the selected positive decision rate has a substantial impact on this cost, an aspect that is not taken into account in previous evaluations.
@@ 2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$. Let $\mathrm{H}_{\mathcal{D}}[c_{w}]$be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\mathcal{D}$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
$\mathrm{H}_{\mathcal{D}}[c_{w}]=\mathrm{E}_{x\in\mathcal{D}}H(y_{x},c_{w}(x))$ (1)
Such measures of a harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases that go unflagged for follow-up treatment).
In fair classification, we usually want to measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate (Verma and Rubin, 2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic loss) while ensuring that the harm is the same per group.
This means that we are searching for a solution to the following problem (where $\mathcal{G}$ is a partitioning of the distribution into groups with respect to a particular protected attribute):
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }$ (2)
$\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\mathcal{G}$
Typical notions of fairness include demographic parity, where the harm corresponds to $1-$ the selection rate, and equal opportunity where it corresponds to the false negative rate.
@@ 2.1 Leveling Down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in a higher false negative rate for some groups i.e. they receiving a lower rate of cancer detection than they would otherwise). This process of enforcing fairness can alter the overall selection rate of the model (Goethals et al., 2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $k$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }\mathrm{H}_{g}[c_{w}]\leq k\,\forall g\in\mathcal{G}$ (3)
This is the same as saying that in the worst case, the harm should be below k, for any group. The challenge now becomes, “How should $k$ be set?” When deciding to deploy a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can be used to select $k$ instances per group.
@@
There are multiple possible answers here. One is a matter of personal utility, that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy, (Menon and Williamson, 2018) and fairness on the ground, which measures the effect of the actual implementation of fairness principles in real-world situations (Bakalar et al., 2021). Constrained resources provide another answer. Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, real-world scenarios often lack the necessary resources to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from a testing to production environment.
@@ 2.2 Constrained Resources
We formalize the notion of resources, as the number of instances predicted as positive by the machine learning model. This term can be used interchangeably with capacity, or with selection rate, positive decision rate or positive prediction rate when speak about the proportion of positively predicted instances. We define $\mathrm{r}_{\mathcal{D}}[c_{w}]$to be the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\mathcal{D}$.
@@
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\text{ such that }\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$ (4)
where $\mathrm{r}_{\mathcal{D}}[c_{w}]$is a weighted sum of group-specific rates $\mathrm{r}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{r}_{\mathcal{D}}[c_{w}]$strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$.
@@
We define $W$ as the set of worst-off groups such that:
@@
The optimum must occur when $\mathrm{r}_{\mathcal{D}}[c_{w}]=r$. If not, one could distribute the remaining resources such that the rates $\mathrm{r}_{g}[c_{w}]$ for the worst-off groups increase without exceeding r, reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{r}_{g}[c_{w}]$).
Furthermore, the optimum is reached when the harm levels are equal across all groups (so all groups are in $W$). If this were not the case, then there would be at least one group not in $W$, that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{r}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$.
The proof also holds when the harm considered is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$ (e.g., a harm corresponding to $1-\text{precision}$). Here, the optimal solution is trivial with an overall selection rate of 0, and equality will be preserved. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some r. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. Revisiting the previous proof, we find that equality holds.
Hence, we find that equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below some value). ∎
@@ 2.3 Cost of Fairness
With this in mind, we consider the following question: What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
Under constrained resources, this leads to the same results as enforcing fairness, i.e. equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
$\mathrm{H}_{\mathcal{D}}[c^{f}_{w}]-\mathrm{H}_{\mathcal{D}}[c^{o}_{w}]$
@@
We might also be interested in performance metrics outside the fairness metric enforced. To quantify the loss in precision, we ask: What is the change in precision if we minimize the harm of the worst-off group? This represents a common scenario. In a medical context it corresponds to the question: What proportion of healthy patients instead of sick will we see if we increase test sensitivity for the worst off groups? In the results, we measure this cost and its contributory factors. This leads to a follow-up question: What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we minimise the harm of the worst off-group? The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increases the resources available, and target these resources at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources. The number of available resources can be an alterable parameter that should explicitly be taken into account.
3 Resource-constrained Fairness and its Cost
While standard fairness methods do not enforce fairness as the selection rate varies (see Figure 2), there is a straightforward algorithm for enforcing fairness at any selection rate, based on the harms based analysis of section 2.2. Given an existing classifier $c_{w}$, we say that a classifier makes a positive decision regarding datapoint $x$ if $c_{w}(x)>t$, where t is some threshold. We consider fairness as in Equation LABEL:eq:fair_class where some notion of harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$is being equalized across groups.
Now, we simply vary a per group threshold $t_{g}$, while computing the group harm $\mathrm{H}_{g}[c_{w}]$, and the corresponding selection rate $\mathrm{r}_{g}[c_{w}]$. We consider some $h$, and select the set of thresholds so that the harm per group lies just under this $h$. This is (approximately) fair with respect to harm $H$, and has a global selection rate
$r=\frac{\sum_{g}|g|\,\mathrm{r}_{g}[c_{w}>t_{g}]}{\sum_{g}|g|}$ (6)
where $|g|$ is the number of datapoints in group $g$. By sweeping over possible values $h$, and examining the corresponding global selection rate, we can find a fair solution that lies just under a target selection rate. The corresponding thresholds can be used to enforce fairness on previously unseen test data.
The resource allocation process is particularly straightforward for demographic parity. In this case the harm corresponds to 1- the selection rate, and the process simplifies to just selecting the top $r$ from each group.  [^4] Note that in particular case of demographic parity ideal thresholds can be calculated on the test set, but for most other measures the thresholds need to be calculated on a separate validation set, as they require access to the target label in order to compute the harms. Owing to sampling error, thresholds selected on the validation data do not perfectly correspond to selection rates on unseen data, and to ensure that the number of instances predicted as positive exactly matches the available resource level, we do not directly use the thresholds. Instead, we calculate the proportion of resources to allocate to each group for each resource level $R$ on the validation set, and then enforce this on the test set.
Our strategy belongs to the class of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained (Mehrabi et al., 2021). We conduct this analysis across a complete range of selection rates $r$ (from 1% to 100 %) to represent different resource levels.
Given this approach, we can now quantify the cost of fairness, defined as the difference in performance (measured as either precision or recall) between the default and fair labeling approaches, for a fixed number of selected instances (R). We vary the available resources $R$ (and consequently the selection rate $r$) and investigate how this influences the cost of fairness.
3.1 Datasets and Classifiers
We use a standard train-test split, where the model is trained on the training set and the predictive performance and cost of fairness is evaluated on a separate test set.
@@ 3.1 Datasets and Classifiers
We use several real-world datasets common in tabular fair machine learning (Le Quy et al., 2022). We also extend our findings to two datasets (CelebA Liu et al. (2015) and Fitzpatrick17k Groh et al. (2021)) from computer vision. The Adult Income dataset is taken from from the 1994 census data, and contains a prediction task identifying if an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. (2020) and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We prepossess the data and binarize the labels following the approach of Zong et al. (2022). We specify the protected and target attributes in Table 1. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $ns$ representing the protected group and $s$ the advantaged group it can be mathematically expressed as:
@@
Classifiers
For the tabular datasets, we use eXtreme Gradient Boosting (XGBoost) Chen and Guestrin (2016). We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. (2016) pretrained on ImageNet Deng et al. (2009) for feature extraction.